OpenStudy (anonymous):

500 grams of Murphonium is found to decay radioactively to 200g in 5 hours. If this decay is exponential, what is the half-life of Murphonium? Note: Not a real substance or science problem, also use the half life equation to solve (create an equation, then solve).

OpenStudy (anonymous):

man, there is a half life equation already in your textbook

OpenStudy (anonymous):

I don't have one...

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

\[f(t) = b * (\frac{ 1 }{ 2 })^{\frac{ t }{ h }}\]

OpenStudy (anonymous):

b = 500 grams t = 200 grams h = 5

OpenStudy (anonymous):

I got 0.0000000045

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

That doesn't make sense

OpenStudy (anonymous):

well that's the equation in my book

OpenStudy (anonymous):

well what are you looking for

OpenStudy (anonymous):

wait sorrry

OpenStudy (anonymous):

do this b = 500 f(t) = 200 t = 5

OpenStudy (anonymous):

now find h

OpenStudy (anonymous):

h is half-life

OpenStudy (anonymous):

\[200 = 500*(\frac{ 1 }{ 2 })^{\frac{ 5 }{ h}}\]

OpenStudy (anonymous):

now find h

OpenStudy (anonymous):

@CaptainSwaggy

OpenStudy (anonymous):

now you can solve this equation using properties of logs to find your half-life @CaptainSwaggy

OpenStudy (anonymous):

@CaptainSwaggy @CaptainSwaggy @CaptainSwaggy @CaptainSwaggy

OpenStudy (anonymous):

I thought you have me H